Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 146     Entry time: 11 Aug 2024, 23:07
Author: Raphael de Wijn 
Group: SPB 
Subject: Shift summary 

X-ray delivery

11.56 keV 202 pulses @0.5 MHz, ~1.0 mJ

some downtime due to fire alarm in the tunnel

Optical laser delivery

800 ns laser for jet visualization

PPLaser for time-resolved data collection

Achievements/Observations

  • PPLaser timing and alingment done
  • Collected several hours of PYP sample with 30ps laser delay

Issues

  • DAMNIT issue solved with OCD DA
  • some online calibration issues from before persisted
ELOG V3.1.4-7c3fd00